# Adaptive Gradient Normalization and Independent Sampling for (Stochastic) Generalized-Smooth Optimization

### v2

@@ Appendix D Proof of Descent Lemma under Generalized PŁ condition
Lemma 5 in (Chen et al., 2023). For any $x\geq 0$, $C\in[0,1]$, $\Delta>0$, and $0\leq w\leq w^{\prime}$ such that $\Delta\geq w^{\prime}-w$, we have the following inequality hold
@@
Let Assumption 1 and 2 hold. Apply AN-GD with $\beta\in[\alpha,1]$. Choose target error $0<\epsilon\leq\min\{1/2\mu\}$. Let step size $\gamma=\frac{(2\mu\epsilon)^{\beta/\rho}}{8(L_{0}+L_{1})+1}$. Denote $\Delta_{t}=f(w_{t})-f^{*}$, then we have
@@
where (i) follows from the descent lemma 1; (ii) follows from the update rule of AN-GD by replacing $w_{t+1}-w_{t}$ with $-\gamma\frac{\nabla f(w_{t})}{\|\nabla f(w_{t})\|^{\beta}}$; (iii) follows from the above technical lemma 5 in (Chen et al., 2023) by letting $\omega^{\prime}=2-\beta$, $\Delta=\beta$ and substitute equation 29 to $2L_{0}\gamma\|\nabla f(w_{t})\|^{2-2\beta}$ and $2L_{1}\gamma\|\nabla f(w_{t})\|^{2+\alpha-2\beta}$ (Let $x=\|\nabla f(w_{t})\|$ and $C=2L_{0}\gamma,2L_{1}\gamma$ , $C\in[0,1]$ because $\epsilon\leq 1/2\mu$); (iv) follows from the fact that $a^{\tau}+b^{\tau}\leq(a+b)^{\tau}$ with $\tau={2}/{\beta}-1\geq 1$ and $a,b\geq 0$, $(v)$ follows from the step size rule $\gamma={(2\mu\epsilon)^{\beta/\rho}}/{(8(L_{0}+L_{1})+1)}$; (vi) follows from the facts that $12L_{0}+2L_{1}\leq 8(L_{0}+L_{1})+$ 1 and $(2L_{0}+2L_{1})^{\frac{2}{\beta}-1}\leq(8(L_{0}+L_{1})+1)^{\frac{2}{\beta}-1}$ when $\beta>0$.
@@
under the step size rule $\gamma=\frac{(2\mu\epsilon)^{\beta/\rho}}{8(L_{0}+L_{1})+1}$.
Moreover, by the generalized PŁ-condition in Definition 5, we have
@@ Appendix E Proof of Theorem 1
$C=\min\Big{\{}\frac{\gamma(2\mu)^{\frac{2-\beta}{\rho}}}{8},\frac{\rho}{2-\beta-\rho}(2^{\frac{2-\beta-\rho}{2-\beta}}-1)\Delta_{0}^{\frac{2-\beta-\rho}{\rho}}\Big{\}}.$
@@
In order to obtain $\Delta_{T}\leq\epsilon$, when $C=\frac{\rho}{2-\beta-\rho}(2^{{(2-\beta-\rho)}/{(2-\beta)}}-1)\Delta_{0}^{{(2-\beta-\rho)}/{\rho}}$, T should satisfy
$T=\frac{1}{(2^{(2-\beta-\rho)/(2-\beta)}-1)(\Delta_{0}\epsilon)^{(2-\beta-\rho)/{\rho}}};$ (39)
@@
$T\geq\max\Big{\{}\frac{8\rho(8(L_{0}+L_{1})+1)}{(2-\beta-\rho)(2\mu)^{2/\rho}\epsilon^{(2-\rho)/{\rho}}},\frac{1}{(2^{(2-\beta-\rho)/(2-\beta)}-1)(\Delta_{0}\epsilon)^{(2-\beta-\rho)/{\rho}}}\Big{\}}=\Omega\big{(}(\frac{1}{\epsilon})^{\frac{2-\rho}{\rho}}\big{)}.$ (41)
@@
To simplify analysis, denote all parameters associated with $\epsilon$ as $\hat{C}=(\frac{C(2\mu)^{{\beta}/{\rho}}}{8(L_{0}+L_{1})+1})^{{\omega}/{\omega-1}}$. We have $(C\gamma)^{{\omega}/{\omega-1}}=\hat{C}\epsilon^{\beta/{\rho+\beta-2}}\leq\hat{C}$. This enables us to further reduce the recursion to
@@
$t-T_{0}$
$\geq$ $\frac{1}{\log\big{(}\frac{\rho}{2-\beta})}\cdot\Big{[}\log\Big{(}\frac{\beta}{\rho+\beta-2}\cdot\log\Big{(}\frac{1}{\epsilon}\Big{)}+\log\Big{(}\log\Big{(}\frac{(2\mu)^{(2+\beta)/\rho}}{4(8(L_{0}+L_{1})+1)^{2}\epsilon}\Big{)}\Big{)}+\log\Big{(}\frac{(4(8(L_{0}+L_{1})+1)^{2})^{{\rho}/{\rho+\beta-2}}\Delta_{T_{0}}}{(2\mu)^{(2+\beta)/(\rho+\beta-2)}}\Big{)}\Big{]}$
$\gtrsim$ $\Omega\Big{(}\log\Big{(}\frac{\beta}{\rho+\beta-2}\cdot\log\Big{(}\frac{1}{\epsilon}\Big{)}+\log\Big{(}\log\Big{(}\frac{(2\mu)^{(2+\beta)/\rho}}{4(8(L_{0}+L_{1})+1)^{2}\epsilon}\Big{)}\Big{)}\Big{)}=\Omega\big{(}\log\big{(}(\frac{1}{\epsilon})^{\frac{\beta}{\rho+\beta-2}}\big{)}\big{)}.$
@@ Appendix H Lemma 5 and Proof
where the first inequality uses $\|\nabla F(w_{t})\|^{\alpha}\leq(1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1}))^{\alpha/2}$ and $(1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)})^{\alpha/2}\geq 1$; the second inequality uses the fact that $\frac{1}{h_{t}}\leq 1$ (so does $\frac{1}{h_{t}^{2\beta}}$) and upper bound ${4\tau_{2}^{2}}$ by $1+2{\tau_{2}^{2}}$; the last inequality uses the fact that thus $(1+4{\tau_{2}^{2}})^{1+\alpha/2}\leq(1+4{\tau_{2}^{2}})^{2}$.
When $\big{\|}\nabla F(w_{t})\big{\|}\geq\sqrt{1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)}}$, we have
@@
$\overset{(ii)}{\leq}$ $\frac{\gamma^{2}}{2h_{t}^{\beta}}\frac{L_{1}\big{\|}\nabla F(w_{t})\big{\|}^{\alpha}}{(4L_{1}\gamma)(\frac{1}{1-\tau_{1}}\big{\|}\nabla f_{\xi^{\prime}}(w_{t})\big{\|}+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\cdot\big{\|}\nabla F(w_{t})\big{\|}^{2}$
@@
$\overset{(iv)}{\leq}$ $\frac{\gamma^{2}}{2h_{t}^{\beta}}\frac{L_{1}}{4L_{1}\gamma}\big{\|}\nabla F(w_{t})\big{\|}^{2}$
@@
where (i) uses the fact $\frac{1}{h_{t}^{\beta}}\leq\frac{1}{4L_{1}\gamma(2\tau_{1}^{2}+1)(\frac{1}{1-\tau_{1}}\|\nabla f_{\xi^{\prime}}(w_{t})\|+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}$; (ii) uses the fact that now $\|\nabla F(w_{t})\|^{2}\geq 1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)}\geq{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)}$; (iii) uses the fact $\|\nabla F(w_{t})\|\leq\frac{1}{1-\tau_{1}}\|\nabla f_{\xi^{\prime}}(w_{t})\|+{\tau_{2}}/{(1-\tau_{1})}$; (iv) uses the fact $\big{\|}\nabla F(w_{t})\big{\|}\geq\sqrt{1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)}}\geq 1$ and thus $\frac{1}{\|\nabla F(w_{t})\|^{\beta-\alpha}}<1$.
Similarly, when $\big{\|}\nabla F(w_{t})\big{\|}\geq\sqrt{1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1})}$, for $\frac{1}{2h_{t}^{2\beta}}\gamma^{2}L_{0}\cdot 4\tau_{2}^{2}$, we have
@@
$\leq$ $\frac{\gamma}{8h_{t}^{\beta}}\big{\|}\nabla F(w_{t})\big{\|}^{2},$ (51)
where the first inequality uses the fact that $\frac{1}{h_{t}^{\beta}}\leq 1$; the second inequality uses the fact $\gamma\leq\frac{1}{4L_{0}(2\tau_{1}^{2}+1)}$; the third inequality uses the fact $\frac{4\tau_{2}^{2}}{2\tau_{1}^{2}+1}\leq 1+\frac{4\tau_{2}^{2}}{2\tau_{1}^{2}+1}$; the forth inequality uses the fact that $\big{\|}\nabla F(w_{t})\big{\|}^{2}\geq 1+{4\tau_{2}^{2}}/{(\tau_{1}^{2}}+\frac{1}{2})$.
Combining equation 49, equation 50, equation 51 gives us the desired result.
@@ Appendix I Proof of Theorem 2
Next, we provide an upper bound of the term $\gamma(2\tau_{1}^{2}+1)\cdot\frac{L_{0}+L_{1}\|\nabla F(w_{t})\|^{\alpha}}{h_{t}^{\beta}}$ in the above equation. By the clipping structure and step size rule, we have $\frac{1}{h_{t}^{\beta}}=\min\big{\{}1,\frac{1}{4L_{1}(2\tau_{1}^{2}+1)\gamma(\frac{1}{1-\tau_{1}}\|\nabla f_{\xi^{\prime}}(w_{t})\|+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\big{\}}<1$ and $\gamma\leq\frac{1}{4L_{0}(2\tau_{1}^{2}+1)}$, which imply that
@@
where (i) utilizes the fact $h_{t}\geq 1$ and $\beta\geq\alpha$; (ii) utilizes the fact that $h_{t}^{\beta}\geq 4L_{1}(2\tau_{1}^{2}+1)\gamma(\frac{1}{1-\tau_{1}}\big{\|}\nabla f_{\xi^{\prime}}(w_{t})\big{\|}+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}$; (iii) utilizes the fact that $\gamma\leq\frac{1}{4L_{1}(2\tau_{1}^{2}+1)}$, thus $(4\gamma L_{1}(2\tau_{1}^{2}+1))\leq(4\gamma L_{1}(2\tau_{1}^{2}+1))^{\alpha/\beta}$ since $\beta\in[\alpha,1]$; (iv) follows from Lemma 2.
@@
where (i) expands the expression of $\frac{1}{h_{t}^{\beta}}$; (ii) utilizes the second inequality of Lemma 2 to upper bound $\|\nabla f_{\xi^{\prime}}(w_{t})\|$ by $(\tau_{1}+1)\|\nabla F(w_{t})\|+\tau_{2}$; (iii) utilizes the fact $\frac{1}{(a+b)^{\beta}}\geq\min\{\frac{1}{(2a)^{\beta}},\frac{1}{(2b)^{\beta}}\}$ by setting $a=\frac{\tau_{1}+1}{1-\tau_{1}}\big{\|}\nabla F(w_{t})\big{\|}$, $b=\frac{2\tau_{2}}{1-\tau_{1}}$, $\beta\geq 0$; (iv) substitute $\gamma$ inside the minimum operator. Moreover, (v) follows from the step size rule $\gamma\leq\frac{1}{8L_{1}(2\tau_{1}^{2}+1)(4\tau_{2}/(1-\tau_{1}))^{\beta}}$. Substituting equation 59 into equation 58 and summing over $t$ from 00 to $T-1$, we obtain that
@@
$\quad\big{(}\Lambda\frac{32L_{1}(2\tau_{1}^{2}+1)((2\tau_{1}+2)/1-\tau_{1})^{\beta}}{T}\big{)}^{\frac{1}{2-\beta}}\Big{\}},$
@@
$\gamma\leq(2\mu\epsilon)^{\frac{4-2\beta}{\rho}}\cdot\frac{1}{4L_{1}(2\tau_{1}+1)/(1-\tau_{1})(2\tau_{1}^{2}+1)}\leq$ $\frac{1}{4L_{1}(2\tau_{1}^{2}+1)}$
@@
where (i)-(v) follows the same arguments stated in Proof G. (vi) applies equation 29 by setting $C=(L_{1}\gamma(2\tau_{1}^{2}+1)^{\beta/2})$ (since $L_{1}\gamma(2\tau_{1}^{2}+1)<\frac{1-\tau_{1}}{2\tau_{2}+2}<1$), $x=\|\nabla F(w_{t})\|$, $\Delta=\beta$, $\omega^{\prime}=2$; (vii) follows from $(L_{1}(2\tau_{1}^{2}+1))^{\beta/2}\geq(L_{1}(2\tau_{1}^{2}+1))^{1/2}$ as $(L_{1}(2\tau_{1}^{2}+1))<1$; (viii) follows from fact for any $b\geq 0$, $\min\{a-b,c\}\geq\min\{a,c\}-b$;(x) follows from fact $\gamma\leq\frac{1}{4L_{1}(2\tau_{1}+2)/(1-\tau_{1})(2\tau_{1}^{2}+1)}$, and thus $\gamma^{3/2}(L_{1}(2\tau_{1}^{2}+1)^{1/2})\geq\frac{1}{4(2C_{1})^{\beta}(2\tau_{1}^{2}+1)}$; (ix) substitutes assumption 2 to replace $\|\nabla F(w_{t})\|$ by $(2\mu\Delta_{t})^{1/\rho}$.
@@
$\mathbb{E}_{w_{t+1}}[\Delta_{t+1}]\leq\mathbb{E}_{w_{t}}[\Delta_{t}-\frac{\gamma^{3/2}(L_{1}(2\tau_{1}^{2}+1))^{1/2}(2\mu\Delta_{t})^{(2-\beta)/\rho}}{4}+\frac{1}{2}\gamma^{2}\big{(}(L_{0}+L_{1})(1+4\tau_{2}^{2})^{2}+{L_{1}(\tau_{1}^{2}+1/2)})]$ (68)

### v3

@@ Appendix D Proof of Descent Lemma under Generalized PŁ condition
Lemma 5 in Chen et al. (2023). For any $x\geq 0$, $C\in[0,1]$, $\Delta>0$, and $0\leq w\leq w^{\prime}$ such that $\Delta\geq w^{\prime}-w$, we have the following inequality hold
@@
Let Assumption 1 and 2 hold. Apply AN-GD with $\beta\in[\alpha,1]$. Choose target error $0<\epsilon\leq\min\{1,1/2\mu\}$. Let step size $\gamma=\frac{(2\mu\epsilon)^{\beta/\rho}}{8(L_{0}+L_{1})+1}$. Denote $\Delta_{t}=f(w_{t})-f^{*}$, then we have
@@
where (i) follows from the descent lemma 1; (ii) follows from the AN-GD update rule by replacing $w_{t+1}-w_{t}$ with $-\gamma\frac{\nabla f(w_{t})}{\|\nabla f(w_{t})\|^{\beta}}$; (iii) follows from the technical lemma 5 above proved in Chen et al. (2023) by letting $\omega^{\prime}=2-\beta$, $\Delta=\beta$ and substituting equation 29 to $2L_{0}\gamma\|\nabla f(w_{t})\|^{2-2\beta}$ and $2L_{1}\gamma\|\nabla f(w_{t})\|^{2+\alpha-2\beta}$ (Let $x=\|\nabla f(w_{t})\|$ and $C=2L_{0}\gamma\text{ or }2L_{1}\gamma$ , $C\in[0,1]$ because $\epsilon\leq 1/2\mu$); (iv) follows from the fact that $a^{\tau}+b^{\tau}\leq(a+b)^{\tau}$ with $\tau={2}/{\beta}-1\geq 1$ and $a,b\geq 0$, $(v)$ follows from the step size rule $\gamma={(2\mu\epsilon)^{\beta/\rho}}/{(8(L_{0}+L_{1})+1)}$; (vi) follows from the facts that $12L_{0}+2L_{1}\leq 8(L_{0}+L_{1})+$ 1 and $(2L_{0}+2L_{1})^{\frac{2}{\beta}-1}\leq(8(L_{0}+L_{1})+1)^{\frac{2}{\beta}-1}$ when $\beta>0$.
@@
under the step size rule $\gamma=\frac{(2\mu\epsilon)^{\beta/\rho}}{8(L_{0}+L_{1})+1}$. Moreover, by the generalized PŁ-condition in Definition 5, we have
@@ Appendix E Proof of Theorem 1
$C=\min\Big{\{}\frac{\gamma(2\mu)^{\frac{2-\beta}{\rho}}}{8},\frac{\rho}{2-\beta-\rho}(2^{\frac{2-\beta-\rho}{2-\beta}}-1)\Delta_{0}^{\frac{\beta+\rho-2}{\rho}}\Big{\}}.$
@@
In order to obtain $\Delta_{T}\leq\epsilon$, when $C=\frac{\rho}{2-\beta-\rho}(2^{{(2-\beta-\rho)}/{(2-\beta)}}-1)\Delta_{0}^{{(\beta+\rho-2)}/{\rho}}$, T should satisfy
$T=\frac{1}{(2^{(2-\beta-\rho)/(2-\beta)}-1)\Delta_{0}^{{(\rho+\beta-2)}/{\rho}}\epsilon^{(2-\beta-\rho)/{\rho}}};$ (39)
@@
$T\geq\max\Big{\{}\frac{8\rho(8(L_{0}+L_{1})+1)}{(2-\beta-\rho)(2\mu)^{2/\rho}\epsilon^{(2-\rho)/{\rho}}},\frac{1}{(2^{(2-\beta-\rho)/(2-\beta)}-1)\Delta_{0}^{{(\rho+\beta-2)}/{\rho}}\epsilon^{(2-\beta-\rho)/{\rho}}}\Big{\}}=\Omega\big{(}(\frac{1}{\epsilon})^{\frac{2-\rho}{\rho}}\big{)}.$ (41)
@@
To simplify analysis, denote all parameters associated with $\epsilon$ as $\hat{C}=(\frac{C(2\mu)^{{\beta}/{\rho}}}{8(L_{0}+L_{1})+1})^{{\omega}/{\omega-1}}$. We have $(C\gamma)^{{\omega}/{\omega-1}}=\hat{C}\epsilon^{\beta/{(\rho+\beta-2)}}\leq\hat{C}$. This enables us to further reduce the recursion to
@@
$t-T_{0}\geq$ $\frac{1}{\log\big{(}\frac{\rho}{2-\beta})}\cdot\Big{[}\log\Big{(}\frac{\beta}{\rho+\beta-2}\cdot\log\Big{(}\frac{1}{\epsilon}\Big{)}+\log\Big{(}\log\Big{(}\frac{(2\mu)^{2/(\rho+\beta-2)}}{(32(L_{0}+L_{1})+4)^{{\rho}/{(\rho+\beta-2)}}\epsilon}\Big{)}\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\log\Big{(}\frac{(32(L_{0}+L_{1})+4)^{{\rho}/{(\rho+\beta-2)}}\Delta_{T_{0}}}{(2\mu)^{2/(\rho+\beta-2)}}\Big{)}\Big{]}$
$\gtrsim$ $\Omega\Big{(}\log\Big{(}\frac{\beta}{\rho+\beta-2}\cdot\log\Big{(}\frac{1}{\epsilon}\Big{)}+\log\Big{(}\log\Big{(}\frac{(2\mu)^{2/(\rho+\beta-2)}}{(32(L_{0}+L_{1})+4)^{\rho/(\rho+\beta-2)}\epsilon}\Big{)}\Big{)}\Big{)}$
$=$ $\Omega\big{(}\log\big{(}(\frac{1}{\epsilon})^{\frac{\beta}{\rho+\beta-2}}\big{)}\big{)}.$
@@ Appendix H Lemma 5 and Proof
where the first inequality uses $\|\nabla F(w_{t})\|^{\alpha}\leq(1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1}))^{\alpha/2}$ and $(1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)})^{\alpha/2}\geq 1$; the second inequality uses the fact that $\frac{1}{h_{t}}\leq 1$ (so does $\frac{1}{h_{t}^{2\beta}}$) and upper bound ${4\tau_{2}^{2}}$ by $1+4{\tau_{2}^{2}}$; the last inequality uses the fact that thus $(1+4{\tau_{2}^{2}})^{1+\alpha/2}\leq(1+4{\tau_{2}^{2}})^{2}$.
When $\big{\|}\nabla F(w_{t})\big{\|}>\sqrt{1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)}}$, we have
@@
<$\overset{(ii)}{<}$ $\frac{\gamma^{2}}{2h_{t}^{\beta}}\frac{L_{1}\big{\|}\nabla F(w_{t})\big{\|}^{\alpha}}{(4L_{1}\gamma)(\frac{1}{1-\tau_{1}}\big{\|}\nabla f_{\xi^{\prime}}(w_{t})\big{\|}+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\cdot\big{\|}\nabla F(w_{t})\big{\|}^{2}$
@@
<$\overset{(iv)}{<}$ $\frac{\gamma^{2}}{2h_{t}^{\beta}}\frac{L_{1}}{4L_{1}\gamma}\big{\|}\nabla F(w_{t})\big{\|}^{2}$
@@
where (i) uses the fact $\frac{1}{h_{t}^{\beta}}\leq\frac{1}{4L_{1}\gamma(2\tau_{1}^{2}+1)(\frac{1}{1-\tau_{1}}\|\nabla f_{\xi^{\prime}}(w_{t})\|+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}$; (ii) uses the fact that now $\|\nabla F(w_{t})\|^{2}>1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)}\geq{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)}$; (iii) uses the fact $\|\nabla F(w_{t})\|\leq\frac{1}{1-\tau_{1}}\|\nabla f_{\xi^{\prime}}(w_{t})\|+\frac{\tau_{2}}{1-\tau_{1}}$; (iv) uses the fact $\big{\|}\nabla F(w_{t})\big{\|}>\sqrt{1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1)}}\geq 1$ and thus $\frac{1}{\|\nabla F(w_{t})\|^{\beta-\alpha}}<1$.
Similarly, when $\big{\|}\nabla F(w_{t})\big{\|}>\sqrt{1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}+1})}$, for $\frac{1}{2h_{t}^{2\beta}}\gamma^{2}L_{0}\cdot 4\tau_{2}^{2}$, we have
@@
$<$ $\frac{\gamma}{8h_{t}^{\beta}}\big{\|}\nabla F(w_{t})\big{\|}^{2},$ (51)
where the first inequality uses the fact that $\frac{1}{h_{t}^{\beta}}\leq 1$; the second inequality uses the fact $\gamma\leq\frac{1}{4L_{0}(2\tau_{1}^{2}+1)}$; the third inequality uses the fact $\frac{4\tau_{2}^{2}}{2\tau_{1}^{2}+1}\leq 1+\frac{4\tau_{2}^{2}}{2\tau_{1}^{2}+1}$; the forth inequality uses the fact that $\big{\|}\nabla F(w_{t})\big{\|}^{2}>1+{4\tau_{2}^{2}}/{(2\tau_{1}^{2}}+1)$. Combining equation 49, equation 50, equation 51 gives us the desired result.
@@ Appendix I Proof of Theorem 2
Next, we provide an upper bound of the term $\gamma(2\tau_{1}^{2}+1)\cdot\frac{L_{0}+L_{1}\|\nabla F(w_{t})\|^{\alpha}}{h_{t}^{\beta}}$ in the above equation. By the clipping structure and step size rule, we have $\frac{1}{h_{t}^{\beta}}=\min\big{\{}1,\frac{1}{4L_{1}\gamma(2\tau_{1}^{2}+1)(\frac{1}{1-\tau_{1}}\|\nabla f_{\xi^{\prime}}(w_{t})\|+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}}\big{\}}\leq 1$ and $\gamma\leq\frac{1}{4L_{0}(2\tau_{1}^{2}+1)}$, which imply that
@@
where (i) utilizes the fact $h_{t}\geq 1$ and $\beta\geq\alpha$; (ii) utilizes the fact that $h_{t}^{\beta}\geq 4L_{1}\gamma(2\tau_{1}^{2}+1)(\frac{1}{1-\tau_{1}}\big{\|}\nabla f_{\xi^{\prime}}(w_{t})\big{\|}+\frac{\tau_{2}}{1-\tau_{1}})^{\beta}$; (iii) utilizes the fact that $\gamma\leq\frac{1}{4L_{1}(2\tau_{1}^{2}+1)}$, thus $(4\gamma L_{1}(2\tau_{1}^{2}+1))\leq(4\gamma L_{1}(2\tau_{1}^{2}+1))^{\alpha/\beta}$ since $\beta\in[\alpha,1]$; (iv) follows from Lemma 2.
@@
where (i) expands the expression of $\frac{1}{h_{t}^{\beta}}$; (ii) utilizes the second inequality of Lemma 2 to upper bound $\|\nabla f_{\xi^{\prime}}(w_{t})\|$ by $(\tau_{1}+1)\|\nabla F(w_{t})\|+\tau_{2}$; (iii) utilizes the fact $\frac{1}{(a+b)^{\beta}}\geq\min\{\frac{1}{(2a)^{\beta}},\frac{1}{(2b)^{\beta}}\}$ by setting $a=\frac{\tau_{1}+1}{1-\tau_{1}}\big{\|}\nabla F(w_{t})\big{\|}$, $b=\frac{2\tau_{2}}{1-\tau_{1}}$, $\beta\geq 0$; (iv) substitute $\gamma$ inside the minimum operator. Moreover, (v) follows from the step size rule $\gamma=\frac{1}{8L_{1}(2\tau_{1}^{2}+1)(2\tau_{2}/(1-\tau_{1}))^{\beta}}\leq\frac{1}{4L_{1}(2\tau_{1}^{2}+1)(4\tau_{2}/(1-\tau_{1}))^{\beta}}$. Substituting equation 59 into equation 58 and summing over $t$ from 00 to $T-1$, we obtain that
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\quad\big{(}\Lambda\frac{32L_{1}(2\tau_{1}^{2}+1)((2\tau_{1}+2)/1-\tau_{1})^{\beta}}{T}\big{)}^{\frac{1}{2-\beta}}\Big{\}},$
@@
$\gamma\leq(2\mu\epsilon)^{\frac{4-2\beta}{\rho}}\cdot\frac{1}{4L_{1}((2\tau_{1}+2)/(1-\tau_{1}))(2\tau_{1}^{2}+1)}\leq$ $\frac{1}{4L_{1}(2\tau_{1}^{2}+1)}$
@@
where (i)-(v) follows the same arguments stated in Proof G. (vi) applies equation 29 by setting $C=(L_{1}\gamma(2\tau_{1}^{2}+1)^{\beta/2})$ (since $L_{1}\gamma(2\tau_{1}^{2}+1)<1$), $x=\|\nabla F(w_{t})\|$, $\Delta=\beta$, $\omega^{\prime}=2,\omega=2-\beta$; (vii) follows from $(L_{1}\gamma(2\tau_{1}^{2}+1))^{\beta/2}\geq(L_{1}\gamma(2\tau_{1}^{2}+1))^{1/2}$ as $(L_{1}\gamma(2\tau_{1}^{2}+1))<1$; (viii) follows from fact for any $b\geq 0$, $\min\{a-b,c\}\geq\min\{a,c\}-b$;(x) follows from fact $\gamma\leq\frac{1}{4L_{1}((2\tau_{1}+2)/(1-\tau_{1}))(2\tau_{1}^{2}+1)}\leq\frac{1}{4L_{1}(2C_{1})^{\beta}(2\tau_{1}^{2}+1)}$. Thus, $L_{1}\gamma(2\tau_{1}^{2}+1)<1$ and $\gamma^{3/2}(L_{1}(2\tau_{1}^{2}+1))^{1/2}\leq\gamma\leq\frac{1}{4(2C_{1})^{\beta}(2\tau_{1}^{2}+1)}$ always holds; (ix) substitutes assumption 2 to replace $\|\nabla F(w_{t})\|$ by $(2\mu\Delta_{t})^{1/\rho}$.
@@
$\mathbb{E}_{w_{t+1}}[\Delta_{t+1}]\leq\mathbb{E}_{w_{t}}[\Delta_{t}-\frac{\gamma^{3/2}(L_{1}(2\tau_{1}^{2}+1))^{1/2}(2\mu\Delta_{t})^{(2-\beta)/\rho}}{4}+\frac{1}{2}\gamma^{2}\big{(}(L_{0}+L_{1})(1+4\tau_{2}^{2})^{2}+{L_{1}(\tau_{1}^{2}+1/2)})].$
@@
$\mathbb{E}_{w_{t+1}}[\Delta_{t+1}]\leq\mathbb{E}_{w_{t}}[\Delta_{t}-\frac{\gamma^{3/2}(L_{1}(2\tau_{1}^{2}+1))^{1/2}(2\mu\Delta_{t})^{(2-\beta)/\rho}}{4}+\frac{\gamma^{3/2}(L_{1}(2\tau_{1}^{2}+1))^{1/2}(2\mu\epsilon)^{(2-\beta)/\rho}}{8}],$ (68)
which gives the desired result.